# Deep Learning and genetic algorithms for cosmological Bayesian inference speed-up

### v1

@@ VI.3 Case 3
Table 3: Wasserstein distances ramdas2017wasserstein between nested sampling posterior samples without and with neuralike, for each free parameter. The closer the value of this distance is to zero, the more similar are the distributions compared. This distance is implemented in scipy and takes into account the 1D posterior samples and their respective weights. Overall, parameters $\Omega_{m}$, $\Omega_{b}h^{2}$, and $h$ exhibit relatively small distances across all cases. However, in Case 1a, higher values of $w_{0}$ and $w_{a}$ distances are observed due to a higher dlogzstart value and fewer data points used. On the other hand, Case 3 demonstrates smaller (better) distances, attributed to the utilization of more data and a lower dlogzstart value.
@@ Appendix A Genetic algorithms as initial live points
As the first insight into the genetic algorithms to generate the initial live points, we show some results about potential advantages in which genetic algorithms could help a nested sampling execution. In Table 4, we can see some tiny examples in which the use of GA to generate the first live points can reduce computational time without sacrificing the statistical results. However, it is worth noticing that we are using a low number of live points because this is the case in which we observed this advantage, when a higher number of live points is used, in general, NS alone is faster because have points in a sparse region of the search space and the use of GA cluster the points around the optimums losing exploration capacity. Nonetheless, there are possible scenarios in which there could be a low number of live points and in these cases, the incursion of GA to generate the initial sampling points could apport an advantage. This is part of a further study of the exploration in detail of this combination between GA with NS.

### v2

@@ VI.3 Case 3
Table 3: Wasserstein distances ramdas2017wasserstein between nested sampling posterior samples without and with neuralike, for each free parameter. The closer the value of this distance is to zero, the more similar are the distributions compared. This distance is implemented in scipy and takes into account the 1D posterior samples and their respective weights. Overall, parameters $\Omega_{m}$, $\Omega_{b}h^{2}$, and $h$ exhibit relatively small distances across all cases. However, in Case 1a, higher values of $w_{0}$ and $w_{a}$ distances are observed due to a higher dlogz_start value and fewer data points used. On the other hand, Case 3 demonstrates smaller (better) distances, attributed to the utilization of more data and a lower dlogz_start value.
@@ Appendix A Genetic algorithms as initial live points
As the first insight into the genetic algorithms to generate the initial live points, we show some results about potential advantages in which genetic algorithms could help a nested sampling execution. In Table 4, we can see some examples in which the use of GA to generate the first live points can reduce computational time without sacrificing the statistical results. However, it is worth noticing that we are using a low number of live points because this is the case in which we observed this advantage, when a higher number of live points is used, in general, NS alone is faster because have points in a sparse region of the search space and the use of GA cluster the points around the optimums losing exploration capacity. Nonetheless, there are possible scenarios in which there could be a low number of live points and in these cases, the incursion of GA to generate the initial sampling points could apport an advantage. This is part of a further study of the exploration in detail of this combination between GA with NS.